OpenStudy (anonymous):

A scientist counts 35 bacteria present in a culture and finds that the number of bacteria triples each hour. The function y = 35 • 3x models the number of bacteria after x hours. Estimate when there will be about 550 bacteria in the culture.

OpenStudy (pfenn1):

The equation is given\[y=35 \times 3x\]where y is the number of bacteria after x hours. To estimate when there will be about 550 bacteria, solve...\[550=35 \times 3x\]Rearrange and solve for x.

OpenStudy (anonymous):

\[550=35\times3^x\] divide by 35 get \[\frac{110}{7}=3^x\] use change of base get \[\frac{\ln(\frac{110}{7})}{\ln(3)}\] then a calculator

OpenStudy (anonymous):

if it says "triples every hour" then the model is \[35\times 3^x\] not \[35\times 3x\]

OpenStudy (pfenn1):

@satellite73, thanks. I guess I read the equation wrong.

OpenStudy (anonymous):

no you read what was written, but that was wrong. hard to write exponents with no symbol palette

OpenStudy (pfenn1):

Well thanks for clearing that up @satellite73